Your grandmother asks for your help in choosing a certificate of deposit​ (CD) from a bank with a​ one-year maturity and a fixed interest rate. The first certificate of​ deposit, CD​ #1, pays 5.95 percent APR compounded quarterly​, while the second certificate of​ deposit, CD​ #2, pays 6.00 percent APR compounded weekly.

****What is the effective annual rate​ (the EAR) of each​ CD,

If the first certificate of​ deposit, CD​ #1, pays 5.95 percent APR compounded ​, the EAR for the deposit is ____% ​(Round to two decimal​ places.)

***CD​ #2, pays 6.00 percent APR compounded weekly. the EAR for the deposit is ____% ​(Round to two decimal​ places.)

and which CD do you recommend to your​ grandmother?
Answer #1

Solution :-

In case of First COD =

Effective Annual Rate = ( 1 + 0.0595 / 4 )4 - 1

= 1.06084 - 1

= 0.06084

= 6.084%

In case of Second COD =

Effective Annual Rate = ( 1 + 0.06 / 52 )52 - 1

= 1.0618 - 1

= 0.0618

= 6.18%

Choose Second COD as it has higher Effective Annual rate

If there is any doubt please ask in comments

Thank you please rate